# Addressing divergent representations from causal interventions on neural networks

### v2

@@ A.1.3 Visualization
We visualized both original and intervened representations by projecting onto the first two principal components of the covariance matrix of $\mathcal{H}$ and $\hat{\mathcal{H}}$ combined:
$).\text{PCs}=\text{eigenvectors}\big(\text{Cov}(\begin{bmatrix}\mathcal{H}\\ \hat{\mathcal{H}}\end{bmatrix}$$\big)$. (14)
@@
A.1.4 Activation patching in balanced subspaces
@@
A.2 CL Loss Continued
A.2.1 Synthetic Dataset Construction
We constructed a synthetic dataset of intermediate-layer representations $h\in\mathbb{R}^{66}$ with known ground-truth labels $y\in\{1,\dots,10\}$ and two causal feature dimensions, where 66comes from 2 feature dimensions plus 64 concatenated noise dimensions is the total feature dimensionality.
@@ A.2.1 Synthetic Dataset Construction
We augmented each 2D noisy base vector with $64$ independent Gaussian noise features, each sampled from $\mathcal{N}(0,1)$, producing final representations $h\in\mathbb{R}^{2+64}$.
A.2.2 MLP Training
@@
where $B$ is the batch size and $p_{\theta}$ denotes the MLP’s predictive distribution over class labels. Optimization used stochastic gradient descent with learning rate $0.005$ for $1000$ epochs with early stopping using an Adam optimizer (kingma2017adam). The code was implemented in PyTorch.
A.2.3 DAS Training
@@
where $\epsilon_{\text{behavior}}$ and $\epsilon_{\text{CL}}$ are tunable coefficients controlling the strength of each term. We only use values of 0 or 1 for $\epsilon_{\text{behavior}}$ and we explore values of $\epsilon_{\text{CL}}$ referring to it as the CL epsilon in figures. We default to an overall learning rate of $0.05$ and subspace size of $1$ unless otherwise specified. Details of these loss functions are provided in Section 5.
A.2.4 Out-of-Distribution Experimental Details
@@
To perform the OOD CL loss experiments, we partitioned the classes into 2 groups, excluding two classes from the partitions entirely. The groups were chosen so that the Sparse set had strictly greater spacing relative to the Dense set. See a visualization of the Dense and Sparse partitions in Figure 4(c). The trainings consisted of the described DAS training settings specified in Appendix A.2.3 except that the data during training was restricted to the respective data partitions. We show IIA values taken from the test sets of each partition.
Figure 4: Out-of-distribution evaluation for DAS trained on the synthetic task with different loss variants. The dashed lines show results for DAS trained using only the behavioral loss. The grey hue indicates that DAS was trained to classify each of the four grey dots and evaluated on classifying the blue dots into their respective categories, and visa-versa for the blue hue. We see that in this setting, the CL loss provides performance benefits for out-of-distribution evaluation.
A.2.5 Further CL Loss Explorations
@@
In these explorations, we explore DAS learning rate choices, number of mask dimensions, and the number of extra noisy dimensions for the OOD experiments. We show accuracies, EMD divergences, and EMD divergences restricted to the causal dimensions. The EMD values are scaled by the number of extra noisy dimensions concatenated to the input representations. We refer to the EMD measurements along causal dimensions only as the Row EMD. We do this for both the trained partitions and held-out partitions for various DAS trainings.
@@
Figure 5: Within distribution hyperparameter search showing the DAS IIA on validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. All reported values include the behavioral loss with different values of the CL Epsilon (shown on the x-axis).
@@
Figure 6: Out-of-distribution hyperparameter search showing the DAS IIA on validation data for the held-out task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. All reported values include the behavioral loss with different values of the CL Epsilon (shown on the x-axis).
@@
Figure 7: Within distribution hyperparameter search showing the DAS IIA on validation data for the trained task partition. We see the DAS mask dimensions (mask_dims) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. All reported values include the behavioral loss with different values of the CL Epsilon (shown on the x-axis).
@@
Figure 8: OOD hyperparameter search showing the DAS IIA on validation data for the held-out partition. We see the DAS mask dimensions (mask_dims) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. All reported values include the behavioral loss with different values of the CL Epsilon (shown on the x-axis).
Figure 9: OOD hyperparameter search showing the intervened EMD along causal dimensions only on validation data for the held-out partition. We see the DAS mask dimensions (mask_dims) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. All reported values include the behavioral loss with different values of the CL Epsilon (shown on the x-axis).
Figure 10: OOD hyperparameter search showing the intervened EMD along all representational dimensions on validation data for the held-out partition. We see the DAS mask dimensions (mask_dims) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. All reported values include the behavioral loss with different values of the CL Epsilon (shown on the x-axis).

### v3

@@ A.1.3 Visualization
We visualized both original and intervened representations by projecting onto the first two principal components of the covariance matrix of $\mathcal{H}^{\prime}$ and $\hat{\mathcal{H}}$ combined:
$).\text{PCs}=\text{eigenvectors}\big(\text{Cov}(\begin{bmatrix}\mathcal{H}^{\prime}\\ \hat{\mathcal{H}}\end{bmatrix}$$\big)$. (14)
@@
Figure 4: A number of additional divergence measures to demonstrate the difference between the natural and intervened distributions. Each is labeled by its y-axis. Each metric is computed over a random sample of natural vectors to simulate the natural manifold, and a sampled set of intervened or natural vectors for which to measure the distance from the natural distribution. We refer to this distribution as the "compared" distribution. The sampled intervened and natural vectors are always the "ground-truth pair" described at the beginning of Appendix A.1. Nearest Cosine Distance: refers to the cosine distance to the nearest sample in the natural manifold. Multiple sampes in the compared distribution can share the same natural sample. This value is averaged over all compared samples. Nearest L2 Distance refers to the cosine distance to the nearest sample in the natural manifold. Multiple sampes in the compared distribution can share the same natural sample. This value is averaged over all compared samples. Min Cos Pairing refers to the lowest cost pairing where cost is the cosine distance between two samples. Vector pairs are exclusive. This value is normalized by the number of samples. Min L2 Pairing refers to the lowest cost pairing where cost is the L2 distance between two samples. Vector pairs are exclusive. This value is normalized by the number of samples. Local PCA Distance refers to the distance to the manifold created using a local PCA of the nearest neighbors. See Appendix A.1.2. EMD refers to the Earth Mover’s Distance. See Appendix A.1.2. KDE refers to the Kernel Density Estimation score. See Appendix A.1.2. Local Linear Reconstruction refers to the local linear reconstruction error. See Appendix A.1.2.
A.2 Closure under coordinate patching characterizes Cartesian products and hyperrectangles
In this section, we will show that axis-aligned hyperrectangles are the only convex manifold shape that does not have source-target vector pairs that produce off-manifold intervened representations.
For vector dimensions $S\subseteq[d]=\{1,\dots,d\}$ and $h^{\textit{src}},h^{\textit{trg}}\in\mathbb{R}^{d}$, define a coordinate patch
$\mathrm{Patch}_{S}(h^{\textit{src}},h^{\textit{trg}})_{i}\;=\;\begin{cases}h_{i}^{\textit{src}},&i\in S,\\ h_{i}^{\textit{trg}},&i\notin S.\end{cases}$
where $i$ refers to the $i^{\text{th}}$ vector coordinate. A set $\mathcal{M}\subseteq\mathbb{R}^{d}$ is patch-closed if $\mathrm{Patch}_{S}(h^{\textit{src}},h^{\textit{trg}})\in\mathcal{M}$ for all $h^{\textit{src}},h^{\textit{trg}}\in\mathcal{M}$ and all $S\subseteq[d]$.
Let $\pi_{i}:\mathbb{R}^{d}\to\mathbb{R}$ be the $i^{\text{th}}$ coordinate projection and write $I_{i}:=\pi_{i}(\mathcal{M})=\{h_{i}:h\in\mathcal{M}\}$.
Theorem A.2 (Patch-closure $\iff$ product of projections).
Let $\mathcal{M}\subseteq\mathbb{R}^{d}$ be nonempty and let $I_{i}:=\pi_{i}(\mathcal{M})$. Then
 $\mathcal{M}\text{ is patch-closed}\iff\mathcal{M}=I_{1}\times\cdots\times I_{d}.$
Proof. ($\Leftarrow$) Immediate: if $\mathcal{M}=\prod_{i}I_{i}$, then patching replaces coordinates by elements of the corresponding $I_{i}$, so the result stays in $\mathcal{M}$.
($\Rightarrow$) Suppose $\mathcal{M}$ is patch-closed. The inclusion $\mathcal{M}\subseteq\prod_{i}I_{i}$ is tautological. For the reverse inclusion, fix $t=(t_{1},\dots,t_{d})$ with $t_{i}\in I_{i}$. For each $i$ pick $h^{(i)}\in\mathcal{M}$ with $h^{(i)}_{i}=t_{i}$. Define $\hat{h}^{(1)}:=h^{(1)}$ and for $k\geq 2$ set $\hat{h}^{(k)}:=\mathrm{Patch}_{\{k\}}(\hat{h}^{(k-1)},h^{(k)})$. Patch-closure gives $\hat{h}^{(k)}\in\mathcal{M}$, and by construction $\hat{h}^{(k)}_{j}=t_{j}$ for all $j\leq k$. Hence $\hat{h}^{(d)}=t\in\mathcal{M}$, yielding $\prod_{i}I_{i}\subseteq\mathcal{M}$. ∎
Corollary (Convex case $\Rightarrow$ hyperrectangle).
If $\mathcal{M}$ is also convex, then each $I_{i}=\pi_{i}(\mathcal{M})\subset\mathbb{R}$ is convex, hence an interval. Therefore $\mathcal{M}=\prod_{i}I_{i}$ is an axis-aligned hyperrectangle (Cartesian product of intervals). Conversely, any axis-aligned hyperrectangle is patch-closed (and convex).
Implication.
Consequently, any nonempty convex set in $\mathbb{R}^{d}$ that is not an axis-aligned hyperrectangle (e.g., a ball, ellipsoid, or a polytope with non-axis-aligned faces) fails to be patch-closed: there exist $h^{\textit{src}},h^{\textit{trg}}\in\mathcal{M}$ and $S\subseteq[d]$ such that $\mathrm{Patch}_{S}(h^{\textit{src}},h^{\textit{trg}})\notin\mathcal{M}$.
A.2.1 Activation patching in balanced subspaces
@@
A.3 CL loss applied to Boundless DAS
To perform the Boundless DAS experiments, we used the Boundless DAS tutorial provided in the pyvene python package (wu2024pyvene) which reproduces a main result from wu2024alpacadas, and we included the CL loss as a weighted auxiliary objective as described in Section 5.1. The exact task used in this tutorial is one involving continuous valued features, which resulted in few occurrences of valid CL vectors in the provided dataset. In order to obtain exact CL vectors, we generated a token sequence sample that contained a valid CL vector for each intervention sample in the dataset. We left hyperparameter choices the same across all trainings except for the CL loss weight $\epsilon$.
A.4 CL Loss in Synthetic Settings
Here we continue Section 5 with experimental details and additional experiments and results.
A.4.1 Synthetic Dataset Construction
The default synthetic task reported in the results section of Section 5.2 was constructed as a dataset of simulated intermediate-layer representations $h\in\mathbb{R}^{18}$ with known ground-truth labels $y\in\{1,\dots,10\}$ and two causal feature dimensions, where 18comes from 2 feature dimensions plus 16 concatenated noise dimensions is the total feature dimensionality. We split these classes into partition 1 and 2, each consisting of 8 of the 10 classes. The held out classes for partition 1 were contained in partition 2 and visa versa. See Figure 5 for a visualization of the task.
@@ A.4.1 Synthetic Dataset Construction
Figure 5: Visualization of the different synthetic tasks used for Figure 3. The Default Task is split into two partitions, both withholding two classes that are contained in the other partition. The OOD task is also split into two partitions, both consisting of 4 classes. The Dense partition consists of a tighter cluster than the Sparse.
@@
We augmented each 2D noisy base vector with $16$ independent Gaussian noise features, each sampled from $\mathcal{N}(0,1)$, producing final representations $h\in\mathbb{R}^{2+16}$.
A.4.2 MLP Training
@@
where $B$ is the batch size and $p_{\theta}$ denotes the MLP’s predictive distribution over class labels.
We perform the the MLP training on both partitions combined for the default dataset split. Then we perform an alignment function training on each partition independently and test the alignment on the untrained partition. We report the average IIA over both data partitions for DAS analyses over 5 seeds. Note that we use an independent MLP for each partition in the OOD experiment (Appendix A.4.4.
Optimization used stochastic gradient descent with learning rate $0.01$ for $300$ epochs with early stopping using an Adam optimizer (kingma2017adam). The code was implemented in PyTorch.
A.4.3 DAS Training
@@
where $\epsilon_{\text{behavior}}$ and $\epsilon_{\text{CL}}$ are tunable coefficients controlling the strength of each term. We only use values of 0 or 1 for $\epsilon_{\text{behavior}}$ and we explore values of $\epsilon_{\text{CL}}$ referring to it as the CL epsilon in figures. We default to an overall learning rate of $0.01$ and subspace size of $1$ unless otherwise specified. Details of these loss functions are provided in Section 5.
Importantly, we stopped training after loss convergence with a patience of 400 training epochs, and we kept the best DAS alignment matrix decided by IIA, and the best CL alignment matrix by EMD. Furthermore, for these trainings, we used a symmetric invertible linear weight matrix as our alignment function as introduced in grant2025das. Namely, the linear alignment matrix $X$ is constructed as $X=(MM^{\top}+\lambda I)S$ where $M\in R^{d_{m}\times d_{m}}$ is a matrix of learned parameters initially sampled from a centered gaussian distribution with a standard deviation of $\frac{1}{d_{m}}$, $I\in R^{d_{m}\times d_{m}}$ is the identity matrix, $\lambda=0.1$ to prevent singular values equal to 0, and $S\in R^{d_{m}\times d_{m}}$ is a diagonal matrix to learn a sign for each column of $X$ using diagonal values $s_{i,i}=\text{Tanh}(a_{i})+\lambda(\text{sign}(\text{Tanh}(a_{i})))$ where each $a_{i}$ is a learned parameter and $\lambda=0.1$ to prevent 0 values. We perform the the alignment function training on both partitions for each synthetic dataset and test each DAS alignment on the untrained partition. We report the average IIA over both data partitions for DAS analyses.
A.4.4 Out-of-Distribution Experimental Details
@@
To perform the OOD CL loss experiments, we partitioned the classes into 2 non-overlapping groups. Two of the 10 classes were excluded entirely. The groups were chosen so that the Sparse set had strictly greater spacing than the Dense set. See a visualization of the Dense and Sparse partitions in Figure 5. A separate MLP training was performed on each partition individually. Then an alignment function was trained on each partition/classifier tuple using the settings specified in Appendix A.4.3. The alignment functions were then tested on the untrained partition. We report IIA values averaged over the performance on each partition.
A.4.5 Further CL Loss Explorations
@@
In these explorations, we explore DAS learning rate and the number of extra noisy dimensions for the OOD experiments. We show accuracies, EMD divergences, and EMD divergences restricted to the causal dimensions. The EMD values are scaled by the number of extra noisy dimensions. We refer to the EMD measurements along causal dimensions only as the Row EMD. We do this for both the trained partitions and held-out partitions for various DAS trainings.
@@
Figure 6: Out of distribution hyperparameter search showing the DAS IIA on OOD validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.
@@
Figure 7: In distribution hyperparameter search showing the DAS IIA on in-distribution validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.
@@
Figure 8: Out of distribution hyperparameter search showing the DAS row-space EMD on OOD validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.
@@
Figure 9: In distribution hyperparameter search showing the DAS row-space EMD on validation data for the trained task partition. We see the DAS loss learning rate (lr) and extra concatenated noisy input dimensions (extra_dim) across the panel columns and rows. The DAS+CL reported values include the behavioral loss whereas the CL label excludes the behavioral loss. The pink dashed lines represent DAS trained with the behavioral loss only.